# Exploring the Evidence-Based SE Beliefs of Generative AI Tools thanks: 979-8-3315-9147-2/25/\$31.00 ©2025 IEEE

### v4

@@ Abstract
Recent innovations in generative artificial intelligence (AI), primarily powered by large language models (LLMs), have transformed how programmers develop and maintain software. The advanced capabilities of generative AI tools to support software development tasks have led to a rise in their adoption within software engineering (SE) workflows. However, little is known about how AI tools perceive evidence-based beliefs and practices supported by research findings. To this end, we conduct a preliminary evaluation conceptually replicating prior work to explore the “beliefs” of generative AI tools used to support software development tasks. We investigate 17 evidence-based claims posited by empirical SE research across five generative AI tools. Our findings show that generative AI tools have ambiguous beliefs regarding research claims and lack credible evidence to support responses. Based on our results, we provide implications for practitioners integrating generative AI-based systems into development contexts and shed light on future research directions to enhance the reliability and trustworthiness of generative AI—aiming to increase awareness and adoption of evidence-based SE research findings in practice.

### v5

@@ Abstract
Background: Recent innovations in generative artificial intelligence (AI) have transformed how programmers develop and maintain software. The advanced capabilities of generative AI tools in supporting development tasks have led to a rise in their adoption within software engineering (SE) workflows. However, little is known about how AI tools perceive evidence-based practices supported by empirical SE research. Aim: To this end, we explore the “beliefs” of generative AI tools increasingly used to support software development in practice. Method: We conduct a preliminary evaluation conceptually replicating prior work to investigate 17 evidence-based claims across five generative AI tools. Results: Our findings demonstrate generative AI tools have ambiguous beliefs regarding research claims and lack credible evidence to support responses. Conclusions: Based on our results, we provide implications for practitioners integrating generative AI-based systems into development contexts and shed light on future research directions to enhance the reliability and trustworthiness of generative AI—aiming to increase awareness and adoption of evidence-based SE research findings in practice.